OpenStudy (anonymous):

The horizontal distance across your ravine is 100 feet...((Okay how do I do this))?? Written assignment

OpenStudy (anonymous):

It says to use radical functions to figure this out...Where do I begin?

OpenStudy (anonymous):

Looks pretty straightforward to me... just follow the directions buddy and have a great time studying geometry. Looks like you only need the pythagorean theorm!

OpenStudy (anonymous):

I don't know how to do that correctly or how to use radical functions really...can someone please help ! explain this to me... anybody.. I need to turn this in tomorrow..

OpenStudy (anonymous):

Would r =100??? and how am I supposed to find the rest of this stuff for the chart?

OpenStudy (anonymous):

It is quite complicated, but you can do it step by step. 1. No figure 2. Horizontal distance: x; Rise: y; Lengh of the bridge: r Use Pythagor theorem: r^2 = y^2 + x^2 (x = 100) r^2 = y^2 + 10,000.(1) 3) I fill the first line only, tou continue. x x^2 y y^2 rise/run r 100 10,000 2.5 6.25 2.5/100 V (6.25 + 10,000)

OpenStudy (anonymous):

Thank you :) !!! I really appreciate it

OpenStudy (anonymous):

When the slope is 6/100 x = 100; y = 6 ; and r^2 = 36 + 10,000 = 10,036. r = 100.18